OpenStudy (anonymous):

How do I solve log(x^2+5x+16)=1?

OpenStudy (agreene):

does the log have a base, or is it a natural log?

OpenStudy (anonymous):

natural

OpenStudy (anonymous):

then the equation will become x^2+5x+16 = e (e means exponential). So you can solve this equation

OpenStudy (anonymous):

if the base is 10 then the equation will become x^2+5x+16=10 so x=-2 0r x=-3

OpenStudy (agreene):

\[\ln (x^2+5x+16)=1\] \[e ^{\ln x^2+5x+16 } = e ^{1}\] \[x^2+ 5x + 16 = e\] \[ x^2 + 5x = e/16\] I dont know of any solutions to this. I mean, I dont know of any x that will end up giving you e/16...

OpenStudy (anonymous):

Well you could try to shove it up your butt http://www.youtube.com/watch?v=ZGJDnFq6vno